OpenStudy (anonymous):

Find the intervals of concavity and points of inflection of y=(4-x^2)^5 I have already found the first derivative to be -10x(4-x^2)^4 but I am stuck on finding the second derivative . . . Wolfram Alpha keeps saying I am wrong and I have no idea why.

jimthompson5910 (jim_thompson5910):

y=(4-x^2)^5 ---------------------------- y'=5(4-x^2)^4*(-2x) y' = -10x(4-x^2)^4 --------------------------- y'' = -10(4 - x^2)^4 + (-10x)*(4)*(4-x^2)^3*(-2x) y'' = -10(4 - x^2)^4 + 80x^2(4-x^2)^3

OpenStudy (anonymous):

Thanks for some reason I wrote down 2x^2 as the derivative of x^2

jimthompson5910 (jim_thompson5910):

yw